# Generalized Out-of-Distribution Detection: A Survey

### v1

@@ 3 Anomaly Detection & One-Class Novelty Detection: Methodology
Given the homogeneous in-distribution data, a straightforward approach is to estimate the in-distribution density and reject OOD test samples that deviate from the estimated distribution. We summarize density-based methods in Section 3.1. Alternative methods rely on the quality of image reconstruction to distinguish anomalous samples (Section 3.2), or directly learn a decision boundary between ID and OOD data (Section 3.3). We also review distance-based and meta-learning-based methods in Section 3.4 and Section 3.5. Lastly, we conclude with a discussion and present theoretical works in Section 3.6.
@@ 3.1.4 Frequency-based Methods
Previous works also explored frequency domain analysis for anomaly detection. While humans perceive images based on low-frequency components, CNN models can largely depend on high-frequency components for decision making [102, 103]. Methods such as CNN kernel smoothing [102] and spectrum-oriented data augmentation [104] are proposed to suppress the influence of high-frequency components. Other works also find that adversarial attacks on low-frequency components are also difficult to detect, therefore proposing methods to target the phase spectrum [105]. Frequency-based methods mainly focus on sensory AD (especially on detecting adversarial examples), which may not be suitable for semantic AD.
@@ 3.2 Reconstruction-based Methods
The core idea of reconstruction-based methods is that the encoder-decoder framework trained on the ID data usually yields different outcomes for ID and OOD samples. The difference in model performance can be utilized as an indicator for detecting anomalies. The difference of model performance can be measured in the feature space (Section 3.2.1) or by the reconstruction error (Section 3.2.2).
@@ 3.3.3 Self-Supervised Learning
3.4 Distance-based Methods
3.5 Gradient-based Methods
@@ 3.6 Discussion
4 Open Set Recognition & Multi-Class Novelty Detection: Methodology
4.1 Classification-based Methods
4.1.1 EVT-based Uncertainty Calibration
4.1.2 EVT-Free Confidence Enhancement
4.1.3 Unknown Class Generation
4.1.4 Label Space Redesign
Both OSR and multi-class ND focus on the settings when ID contains more than one category. One-hot encoding is commonly used to encode categorical information for classification. However, one-hot encoding ignores the inherent relationship among labels. For example, it is unreasonable to have a uniform distance between dog and cat vs. dog and car. To this end, several works attempt to use information in the label space for novel classes detection. Some works arrange the large semantic space into a hierarchical taxonomy of known classes [167, 168]. Under the redesigned label architecture, top-down classification strategy [167] and group softmax training [168] are demonstrated effective. Another set of works uses word embeddings to automatically construct the label space. In [169], the sparse one-hot labels are replaced with several dense word embeddings from different NLP models, forming multiple regression heads for robust training. When testing, the label, which has the minimal distance to all the embedding vectors from different heads, will be considered as the prediction. If the minimal distance crosses above the threshold, the sample would be classified as “novel”. Recent works further take the image features from language-image pre-training models [170] to better detect novel classes, where the image encoding space also contains rich information from the label space [171, 172].
4.2 Distance-based Methods
Distance-based methods for OSR require the prototypes to be class-conditional, which allows maintaining the ID classification performance. Category-based clustering and prototyping are performed based on the visual features extracted from the classifiers. OOD samples can be detected by computing the distance w.r.t clusters [173, 174]. Some methods also leveraged contrastive learning to learn more compact clusters for known classes [175, 176], which enlarge the distance between ID and OOD. CROSR [177] enhances the features by concatenating visual embeddings from both the classifier and reconstruction model for distance computation in the extended feature space. Besides using features from classifiers, GMVAE [178] extracts features using a reconstruction VAE, and models the embeddings of the training set as a Gaussian mixture with multiple centroids for the following distance-based operations. Classifiers using nearest neighbors are also adapted for OSR problem [179]. By storing the training samples, the nearest neighbor distance ratio is used for identifying unknown samples in testing.
4.3 Reconstruction-based Methods
4.3.1 Sparse Representation Methods
4.3.2 Reconstruction-Error Methods
4.4 Discussion
Due to the restriction on using only ID data for training, OSR methods do not implement background classes, or outlier exposure (more in Section 5.1.1). We proceed by reviewing the recent OOD detection literature, which encompasses a broader problem space and methodological solutions to detecting the unknowns.
Figure 3: Timeline for representative methodologies of (a) anomaly detection & one-class novelty detection, details in Section 3, (b) multi-class novelty detection & open set recognition, details in Section 4, and (c) OOD detection, details in Section 5, under generalized OOD detection framework. Different colors indicate different categories of methodologies. Each method has its corresponding reference (inconspicuous white) in the lower right corner. We do not list outlier detection methods in this figure due to their limited number of works on computer vision in deep learning era.
5 OOD Detection: Methodology
5.1 Classification-based Methods
5.1.1 Output-based Methods
a. Post-hoc Detection  Post-hoc methods have the advantage of being easy to use without modifying the training procedure and objective. The property can be important for the adoption of OOD detection methods in real-world production environments, where the overhead cost of retraining can be prohibitive. Early work ODIN [188] is a post-hoc method that uses temperature scaling and input perturbation to amplify the ID/OOD separability. Key to the method, a sufficiently large temperature has a strong smoothing effect that transforms the softmax score back to the logit space—which effectively distinguishes ID vs. OOD. Note that this is different from confidence calibration, where a much milder $T$ is employed. While calibration focuses on representing the true correctness likelihood of ID data only, the ODIN score is designed to maximize the gap between ID and OOD data and may no longer be meaningful from a predictive confidence standpoint. Built on the insights, recent work [189] proposed using an energy score for OOD detection, which is hyperparameter-free and achieves comparable or even better performance than ODIN. The energy function maps the logit outputs to a scalar through a convenient logsumexp operator. Test samples with lower energy are considered ID and vice versa. [55] further proposed JointEnergy score, which improves OOD uncertainty estimation for multi-label classification networks. Recently, [190] reveals one fundamental cause of the overconfidence issue on OOD data. In particular, using mismatched BatchNorm statistics—that are estimated on ID data yet blindly applied to the OOD data in testing—can trigger abnormally high unit activations and model output accordingly.  [190] provides a simple activation rectification strategy termed ReAct, which establishes strong post-hoc detection performance.
b. Confidence Enhancement Methods  Tailored for OOD detection problem, confidence can be developed via designing a confidence-estimating branch [193] or class [194], data augmentation [195], ensembling with leaving-out strategy [196], adversarial training [197, 198, 199, 200, 214], stronger data augmentation [201, 202, 203, 204], certified certain uncertainty modeling [205], and utilizing feature from the optimal depth [206]. Specially, to enhance the sensitivity to covariate shift, some methods focus on the hidden representations in the middle layers of neural networks. Generalized ODIN, or G-ODIN [58] extended ODIN [188] by using a specialized training objective termed DeConf-C and choose hyperparameters such as perturbation magnitude on ID data. Note that we do not categorize G-ODIN as post-hoc method as it requires model retraining. Techniques such as layer-wise Mahalanobis distance [207] and Gram Matrix [192, 208, 209] are implemented for better hidden feature quality to perform density estimation.
c. Outlier Exposure  Another branch of OOD detection methods makes use of a set of collected OOD samples, or “outlier”, during training to help models learn ID/OOD discrepancy. Starting from the concurrent baselines that encourage a flat/high-entropic prediction on given OOD samples [210, 52] and suppressing OOD feature magnitudes [52], a follow-up work, MCD [211] uses a network with two branches, between which entropy discrepancy is enlarged for OOD training data. Another straightforward approach with outlier exposure spares an extra abstention (or rejection class) and considers all the given OOD samples in this class [219, 214, 217]. A later work OECC [212] noticed that an extra regularization for confidence calibration introduces additional improvement for OE. To effectively utilize the given, usually massive, OOD samples, some works use outlier mining [214] and adversarial resampling [213] approaches to obtain a compact yet representative set. Other works consider a more practical scenario where given OOD samples contain ID samples, therefore using pseudo-labeling [215] or ID filtering methods [216] to reduce the interference of introduced ID. In general, OOD detection with outlier exposure can reach a much better performance. However, as research shows that the performance can be largely affected by the correlations between given and real OOD samples [218], further exploration is still required to generalize the exposed OOD samples to broader, universal OOD samples.
5.1.2 OOD Data Generation
The outlier exposure approaches impose a strong assumption on the availability of OOD training data, which can be infeasible in practice. When no OOD sample is available, some methods attempt to synthesize OOD samples to enable ID/OOD separability. Existing works leverage GANs to generate OOD training samples and force the model predictions to be uniform [220], generate boundary samples in the low-density area [221], or similarly, high-confidence OOD samples [222].
5.1.3 Gradient-based Methods
5.1.4 Bayesian Models
A Bayesian model is a statistical model that implements Bayes’ rule to infer all uncertainty within the model [291]. The most representative method is the Bayesian neural network [292], which draws samples from the posterior distribution of the model via MCMC [293], Laplace methods [294, 295] and variational inference [296], forming the epistemic uncertainty of the model prediction. However, their obvious shortcomings of inaccurate predictions [297] and high computational costs [298] prevent them from wide adoption in practice. Recent works attempt several less principled approximations including MC-dropout [224] and deep ensembles [299, 225] for faster and better estimates of uncertainty. These methods are less competitive for OOD uncertainty estimation. Further exploration takes natural-gradient variational inference and enables practical and affordable modern deep learning training while preserving the benefits of Bayesian principles [226]. Dirichlet Prior Network (DPN) is also used for OOD detection with an uncertainty modeling of three different sources of uncertainty: model uncertainty, data uncertainty, and distributional uncertainty and form a line of works [227, 228, 229].
5.1.5 Large-scale OOD Detection
5.2 Density-based Methods
Density-based methods in OOD detection explicitly model the in-distribution with some probabilistic models and flag test data in low-density regions as OOD. Although OOD detection can be different from AD in that multiple classes exist in the in-distribution, density estimation methods used for AD in Section 3.1.2 can be directly adapted to OOD detection by unifying the ID data as a whole [87, 88, 89, 90, 121]. When the ID contains multiple classes, class-conditional Gaussian distribution can explicitly model the in-distribution so that the OOD samples can be identified based on their likelihoods [207]. Flow-based methods [92, 232, 233, 234] can also be used for probabilistic modeling. While directly estimating the likelihood seems like a natural approach, some works [235, 236, 237] find that probabilistic models sometimes assign a higher likelihood for the OOD sample. Several works attempt to solve the problems using likelihood ratio [238].  [239] finds that the likelihood exhibits a strong bias towards the input complexity and proposes a likelihood ratio-based method to compensate the influence of input complexity. Recent methods turn to new scores such as likelihood regret [240] or an ensemble of multiple density models [236]. Overall, generative models can be prohibitively challenging to train and optimize, and the performance can often lag behind the classification-based approaches (Section 3.3).
5.3 Distance-based Methods
The basic idea of distance-based methods is that the testing OOD samples should be relatively far away from the centroids or prototypes of in-distribution classes.  [207] uses the minimum Mahalanobis distance to all class centroids for detection. A subsequent work splits the images into foreground and background, and then calculates the Mahalanobis distance ratio between the two spaces [241]. Some works use cosine similarity between test sample features and class features to determine OOD samples [242, 243]. The one-dimensional subspace spanned by the first singular vector of the training features is shown to be more suitable for cosine similarity-based detection [244]. Moreover, other works leverage distances with radial basis function kernel [245] and Euclidean distance [246] between the input’s embedding and the class centroids.
5.4 Discussion
The field of OOD detection has enjoyed rapid development since its emergence, with a large space of solutions ranging from classification-based to density-based to distance. In the multi-class setting, the problem can be canonical to OSR (Section 4)—accurately classify test samples from ID within the class space $\mathcal{Y}$, and reject OOD test samples with semantics outside the support of $\mathcal{Y}$. However, OOD detection encompasses a broader spectrum of learning tasks (e.g., multi-label classification [55]) and solution space (e.g., density estimation and outlier exposure). Some approaches relax the constraints imposed by OSR and achieve strong performance.
@@ 6.4 Discussion
Although the application of OD is not as common as other sub-tasks in the computer vision community, techniques for OD can be valuable for other tasks such as open-set semi-supervised learning, learning with open-set noisy labels, and potentially novelty discovery. In addition to methods reviewed here, most solutions to AD/ND/OOD detection can also be applied by considering all observations as ID (for model training) and then applying the model again on all the observations. In this case, methods such as reconstruction-based PCA (c.f. Section 3.2) and energy-based models (c.f. Section 3.1.3 and 5.1.1) can also address the OD problem.
@@ 7.1 Challenges
We also urge the community to carefully construct the benchmarks and evaluations. It is noticed that early work [187] ignored the fact that some OOD datasets may contain images with ID categories, causing inaccurate performance evaluation. Fortunately, recent OOD detection works [168, 216] have realized this flaw and pay special attention to removing ID classes from OOD samples to ensure proper evaluation.
b. Outlier-free OOD Detection  The outlier exposure approach [210] imposes a strong assumption of the availability of OOD training data, which can be difficult to obtain in practice. Moreover, one needs to perform careful de-duplication to ensure that the outlier training data does not contain ID data. These restrictions may lead to inflexible solutions and prevent the adoption of methods in the real world. As with the recent taken-down of TinyImages dataset [310], it poses a reproducibility crisis for OE-based methods. Going forward, a major challenge for the field is to devise outlier-free learning objectives that are less dependent on auxiliary outlier dataset.
@@
d. Real-world Benchmarks and Evaluations  Current methods have been primarily evaluated on small data sets such as CIFAR. It’s been shown that approaches developed on the CIFAR benchmark might not translate effectively into ImageNet benchmark with a large semantic space, highlighting the need to evaluate OOD detection in a large-scale real-world setting. Therefore, we encourage future research to evaluate on ImageNet-based OOD detection benchmark [168], as well as large-scale OSR benchmark [311], and test the limits of the method developed. Moreover, real-world benchmarks that go beyond image classification can be valuable for the research community. In particular, for safety-critical settings such as autonomous driving and medical imaging diagnosis, more specialized benchmarks are needed and should be carefully constructed.
@@ 7.2 Future Directions
b. OOD Detection & Generalization  An open-world classifier should consider two tasks, i.e., being robust to covariate shift while being aware of the semantic shift. Existing works pursue these two goals independently. Recent work proposes a semantically coherent OOD detection framework [216] that encourages detecting semantic OOD samples while being robust to negligible covariate shift. Given the vague definition of OOD, [312] proposed a new formalization of OOD detection by explicitly taking into account the separation between invariant features (semantic related) and environmental features (non-semantic). The work highlighted that spurious environmental features in the training set can significantly impact OOD detection, especially when the label-shifted OOD data contains the spurious feature. Recent works on open long-tailed recognition [80], open compound domain adaptation [76], open-set domain adaptation [313] and open-set domain generalization [314] consider the potential existence of open-class samples. Looking ahead, we envision great research opportunities on how OOD detection and OOD generalization can better enable each other [80], in terms of both algorithmic design and comprehensive performance evaluation.
c. OOD Detection & Open-Set Noisy Labels  Existing methods of learning from open-set noisy labels focus on suppressing the negative effects of noise [68, 315]. However, the open-set noisy samples can be useful for outlier exposure (c.f5.1.1[308] and potentially benefit OOD detection. With a similar idea, the setting of open-set semi-supervised learning can be promising for OOD detection. We believe the combination between OOD detection and the previous two fields can provide more insights and possibilities.
d. Theoretical Analysis  While most of the existing OOD detection works focus on developing effective approaches to obtain better empirical performance, the theoretical analysis remains largely untapped. We hope future research can also contribute theoretical analyses and provide in-depth insights that help guide algorithmic development with rigorous guarantees.

### v2

@@ 5 Anomaly Detection & Novelty Detection: Methodology
Given the homogeneous in-distribution data, a straightforward approach is to estimate the in-distribution density and reject OOD test samples that deviate from the estimated distribution. We summarize density-based methods in Section 5.1. Alternative methods rely on the quality of image reconstruction to distinguish anomalous samples (Section 5.2), or directly learn a decision boundary between ID and OOD data (Section 5.4). We also review distance-based in Section 5.3. Lastly, we conclude with a discussion and present theoretical works in Section 5.5.
@@ 5.1.4 Frequency-based Methods
Previous works also explored frequency domain analysis for anomaly detection. While humans perceive images based on low-frequency components, CNN models can largely depend on high-frequency components for decision-making [227, 228]. Methods such as CNN kernel smoothing [227] and spectrum-oriented data augmentation [229] are proposed to suppress the influence of high-frequency components. Other works also find that adversarial attacks on low-frequency components are also difficult to detect, therefore proposing methods to target the phase spectrum [230]. Frequency-based methods mainly focus on sensory AD (especially on detecting adversarial examples).
@@ 5.2 Reconstruction-based Methods
Similar to Section 4.3, some AD methods rely on the different performance on performing normal and abnormal data to identify anomalies. The difference of model performance can be measured in the feature space (Section 5.2.1) or by the reconstruction error (Section 5.2.2).
@@ 5.2.2 Reconstruction-Error Methods
5.2.3 Gradient-based Methods
5.3 Distance-based Methods
@@ 6.4 Discussion
Although the application of OD is not as common as other sub-tasks in the computer vision community, techniques for OD can be valuable for other tasks such as open-set semi-supervised learning, learning with open-set noisy labels, and potentially novelty discovery. In addition to methods reviewed here, most solutions to AD/ND/OOD detection can also be applied by considering all observations as ID (for model training) and then applying the model again on all the observations. In this case, methods such as reconstruction-based PCA (c.f. Section 5.2) and energy-based models (c.f. Section 5.1.3 and 3.1.1) can also address the OD problem.
@@ 7.1 Challenges
We also urge the community to carefully construct the benchmarks and evaluations. It is noticed that early work [64] ignored the fact that some OOD datasets may contain images with ID categories, causing inaccurate performance evaluation. Fortunately, recent OOD detection works [131] have realized this flaw and pay special attention to removing ID classes from OOD samples to ensure proper evaluation.
b. Outlier-free OOD Detection  The outlier exposure approach [67] imposes a strong assumption of the availability of OOD training data, which can be difficult to obtain in practice. Moreover, one needs to perform careful de-duplication to ensure that the outlier training data does not contain ID data. These restrictions may lead to inflexible solutions and prevent the adoption of methods in the real world. Going forward, a major challenge for the field is to devise outlier-free learning objectives that are less dependent on auxiliary outlier dataset.
@@
d. Real-world Benchmarks and Evaluations  Current methods have been primarily evaluated on small data sets such as CIFAR. It’s been shown that approaches developed on the CIFAR benchmark might not translate effectively into ImageNet benchmark with a large semantic space, highlighting the need to evaluate OOD detection in a large-scale real-world setting. Therefore, we encourage future research to evaluate on ImageNet-based OOD detection benchmark [138], as well as large-scale OSR benchmark [314], and test the limits of the method developed. Moreover, real-world benchmarks that go beyond image classification can be valuable for the research community. In particular, recent works start looking more into object-level OOD detection [146, 147], which can be useful for safety-critical settings such as autonomous driving.
@@ 7.2 Future Directions
b. OOD Detection & Generalization  An open-world classifier should consider two tasks, i.e., being robust to covariate shift while being aware of the semantic shift. Existing works pursue these two goals independently. Recent work proposes a semantically coherent OOD detection framework [131] that encourages detecting semantic OOD samples while being robust to negligible covariate shift. Given the vague definition of OOD, [346] proposed a formalization of OOD detection by explicitly taking into account the separation between invariant features (semantic related) and environmental features (non-semantic). The work highlighted that spurious environmental features in the training set can significantly impact OOD detection, especially when the semantic OOD data contains the spurious feature. Further, full-spectrum OOD detection [308] highlights the effects of “covariate-shifted in-distribution”, and show that most of the previous OOD detectors are unfortunately sensitive to covariate shift rather than semantic shift. This setting explicitly promotes the generalization ability of OOD detectors. Recent works on open long-tailed recognition [100], open compound domain adaptation [92], open-set domain adaptation [347] and open-set domain generalization [348] consider the potential existence of open-class samples. Looking ahead, we envision great research opportunities on how OOD detection and OOD generalization can better enable each other [100], in terms of both algorithmic design and comprehensive performance evaluation.
c. OOD Detection & Open-Set Noisy Labels  Existing methods of learning from open-set noisy labels focus on suppressing the negative effects of noise [84, 349]. However, the open-set noisy samples can be useful for outlier exposure (c.f. Section 3.1.1[345] and potentially benefit OOD detection. With a similar idea, the setting of open-set semi-supervised learning can be promising for OOD detection. We believe the combination between OOD detection and the previous two fields can provide more insights and possibilities.
d. Theoretical Analysis  While most of the existing OOD detection works focus on developing effective approaches to obtain better empirical performance, the theoretical analysis remains largely underexplored. Recently,  [293] developed an analytical framework that provided theoretical understanding for OOD detection. We hope future research can also contribute theoretical analyses and provide in-depth insights that help guide algorithmic development with rigorous guarantees.
e. OOD Detection For Broader Learning Tasks  As mentioned in the Section 3.5, OOD detection encompasses a broader spectrum of learning tasks, including multi-label classification [106], object detection [146, 147], image segmentation [68], time-series prediction [350], and LiDAR-based 3D object detection [351]. For the classification task itself, the researchers also extended the OOD detection technique to improve the reliability of zero-shot pretrained models [352] (e.g., CLIP). Other works explore how to use OOD detection methods to only generate reliable image captions [353]. We believe that with OOD detection methods, many more fields can benefit from reliable and practical models. In turn, insights from various fields can also help OOD detection to develop better.